OpenStudy (anonymous):

Help please. I need to know if I am right to this point and I need help with the rest. Use the given information about a polynomial whose coefficients are real #s to find the remaining zeros of the polynomial. Degree: 6 Zeros: 10+20i, -3-isqrt6, 19-20TTi This is what I have; roots are 10+20i & 10-20i (-3-isqrt6) ( -3 + isqrt6) 19-20TTi & 19+20TTi polynomial is: ?

OpenStudy (anonymous):

9 + 6 = 15 simple as that

OpenStudy (anonymous):

x^2 - y^2 = (x+y)(x-y) i^2 = -1 sqrt of 6 squared = 6 -3^2 = 9

OpenStudy (anonymous):

so it's: -3^2 - i^2(sqrt of 6)

OpenStudy (barrycarter):

Hint: try multiplying the roots in complex conjugate pairs. For example: (x-(10+20I))*(x-(10-20I))